# PerceiverS: A Multi-Scale Perceiver with Effective Segmentation for Long-Term Expressive Symbolic Music Generation

### v2

@@ VI-B2 Experiment 2: Multi-Scale Attention
Results, as shown in the data within the solid blue box in Table I, indicate that this approach significantly improved the model’s generation quality. By examining the OA of the four parameters IOI, IOI/seg, NLH, and NLTM—which are related to rhythm characteristics—it is evident that the model’s generation aligns more closely with the reference set. Additionally, the OA values for PCH and PCH/seg indicate that the harmonic characteristics also better resemble those of the reference dataset. Furthermore, we analyzed the density of repetitive notes detected per unit time in segmented generated music before and after the improvement, as shown in Figure 5, which indicates that the pre-improvement model exhibited a greater tendency to generate repetitive segments. In contrast, unusual repetitive segments in the generated music are now rare after the improvement, as shown in Figure 4, with the generated music even exhibiting a rich diversity of phrases.
@@ VI-C Discussion
Our enhanced model, PerceiverS (Segmentation and Scale), leverages Effective Segmentation aligned with Perceiver AR [4]’s operation mode. By applying Multi-Scale masking strategies within multi-layer cross-attention, PerceiverS effectively generates cohesive, consistent music over extended temporal contexts, preserving intricate musical variations and expressive details across long sequences. Additionally, through the use of performance music datasets, PerceiverS is capable of producing high-quality symbolic music that captures the nuances of human performance. Importantly, because the model does not rely on annotated datasets, PerceiverS can be trained on MIDI data derived from audio using any automatic music transcription (AMT) technique. This capability suggests a future in which symbolic music generation can leverage vast historical recordings, unbounded by the limitations of manually annotated data.
@@ VII Conclusion and Future Work
In this work, we introduced a novel model, PerceiverS, which builds on the Perceiver AR [4] architecture by incorporating Effective Segmentation and a Multi-Scale attention mechanism. The Effective Segmentation approach progressively expands the context segment during training, aligning more closely with auto-regressive generation and enabling smooth, coherent generation across ultra-long symbolic music sequences. The Multi-Scale attention mechanism further enhances the model’s ability to capture long-term structural dependencies while maintaining short-term diversity.

### v3

@@ VI-B2 Experiment 2: Multi-Scale Attention
Results, as shown in the data within the solid blue box in Table I, indicate that this approach significantly improved the model’s generation quality. By examining the OA of the four parameters IOI, IOI/seg, NLH, and NLTM—which are related to rhythm characteristics—it is evident that the model’s generation aligns more closely with the reference set. Additionally, the OA values for PCH and PCH/seg indicate that the harmonic characteristics also better resemble those of the reference dataset. Furthermore, we analyzed the density of repetitive notes detected within each of 64 equal-length segments (per piece), across all generated samples, both before and after the proposed improvements, as shown in Figure 6, which indicates that the pre-improvement model exhibited a greater tendency to generate repetitive segments. In contrast, unusual repetitive segments in the generated music are now rare after the improvement, as shown in Figure 5, with the generated music even exhibiting a rich diversity of phrases.
@@
| Tempo | Motif | Rhythm | Harmony | Style Identified | Tonal | Human Composed | Overall | |
| Coherence | Presence | Irregularity | Progression | Yes (%) | Yes (%) | Yes (%) | Quality | |
| Baseline | 2.86 | 2.14 | 2.71 | 1.64 | 14% | 4% | 18% | 2.00 |
| PerceiverS | 3.82 | 2.82 | 2.00 | 2.79 | 36% | 50% | 29% | 3.29 |
Note: Scores range from 1 (poor) to 5 (excellent). Percentages indicate the proportion of evaluators who selected the affirmative option in each multiple-choice question.
Finally, we invited two professional pianists to evaluate the musical outputs generated by the baseline model and our proposed improved model, PerceiverS. Each model produced 14 pieces for evaluation. The evaluation followed the methodology introduced in the questionnaire from Hernández-Oliván et al.’s work on the Subjective Evaluation of Deep Learning Models for Symbolic Music Composition [29]. It was conducted using a blind listening approach and incorporated both quantitative and qualitative assessments. The questionnaire consisted of a 5-point Likert scale evaluating aspects such as tempo coherence, presence of motifs, rhythmic irregularity, harmonic progression, and overall musical quality. In addition, participants were asked to identify the musical style, assess tonality, and determine whether the piece was composed by a human or generated by AI, following a Turing test format.
The evaluation results are summarized in Table III, where all criteria confirm through human assessment that PerceiverS significantly outperforms the baseline in generated music quality.
@@ VI-C Discussion
Our enhanced model, PerceiverS (Segmentation and Scale), leverages Effective Segmentation aligned with Perceiver AR [4]’s operation mode. By applying Multi-Scale masking strategies within multi-layer cross-attention, PerceiverS effectively generates cohesive, consistent music over extended temporal contexts, preserving intricate musical patterns and expressive details across long sequences. Additionally, through the use of performance music datasets, PerceiverS is capable of producing high-quality symbolic music that captures the nuances of human performance. Importantly, because the model does not rely on annotated datasets, PerceiverS can be trained on MIDI data derived from audio using any automatic music transcription (AMT) technique. This capability suggests a future in which symbolic music generation can leverage vast historical recordings, unbounded by the limitations of manually annotated data.
@@ VII Conclusion and Future Work
In this work, we introduced a novel model, PerceiverS, which builds on the Perceiver AR [4] architecture by incorporating Effective Segmentation and a Multi-Scale attention mechanism. The Effective Segmentation approach progressively expands the context segment during training, aligning more closely with auto-regressive generation and enabling smooth, coherent generation across ultra-long symbolic music sequences. The Multi-Scale attention mechanism further enhances the model’s ability to capture both long-term structural dependencies and short-term expressive details.